Please show me how to solve. Thanks!
5(X-4)-4(x-2)=0

Answer by tutorcecilia(2152)

5(X-4)-4(x-2)=0 [multiply using the distibutive property]
5x-20-4x+8=0 [combine like-terms]
x-12=0 [solve for x]
x=12
.
checking
5(X-4)-4(x-2)=0
5((12)-4)-4((12)-2)=0
5(8)-4(10)=0
40-40=0
0=0 [checks out]